1 The Course

1.1 The website

You can access the website by going to the Department of Statistics website

www.stats.ox.ac.uk/current student/bammath/course material

and finding Part A: Statistics in the Lecture Notes section.

ALTERNATIVELY

Go straight to

www.stats.ox.ac.uk/˜ myers/stats materials.html

The website begins with a STOP PRESS noticeboard which is used for announcements.

There will be sections on:

• Lecture notes: these will be exactly the same as the OHP slides I shall use and will require annotations and additions during the lectures.

• Problem sheets: these are for tutorials - there will be no hard copy hand-outs.

• Data sets used for illustration and practice.

• An introduction to the statistical computer package R. Although the course does not contain compulsory practicals, you may find it useful to try some analyses for yourself. This package is the one used for the third-year course BS1: Applied Statistics and may be downloaded free of charge from the web.

The R package is available free of charge on the web. Go to the web address

www.stats.bris.ac.uk/R/ and download (if you’re using Windows) by following the sequence

Windows (95 and later) → base → R-2.4.1-win32.exe with versions and available similarly for Linux, Mac platforms too. An Introduction to R by W.N. Venables and D.M. Smith can be obtained

by following the sequence Manual → An Introduction to R

1.2 The plan

• ”Skeleton” lecture notes will be provided on the web. You are expected to bring these to lectures and augment them.

• The notes will appear on the website ahead of the lectures in order to give you time to download them and bring them along.

• I shall revise bits of Mods statistics and bits of probability as and when we need them.

• We shall begin the A5 material proper with the important topic of Estimation.

• The next topic will be Hypothesis Testing, which will be entirely new to you.

• Confidence intervals will be revisited in the light of the new theory you have covered.

• The final topic will be Regression, with which you already have a brief acquaintance from the simple treatment you met in Mods.